OpenStudy (anonymous):

The sum of the squares of two consecutive positive even numbers is 20. What is the smaller number? @Noura11

OpenStudy (anonymous):

Let a and b the two consecutive positive even numbers. Then : \[a=2n\text{ and } b=a+2=2n+2\] So we have : \[a^2+b^2=20\] Then : \[4n^2+(2n+2)^2=20\] Then : \[4n^2+4n^2+8n+4=20\] Then : \[8n^2+8n-16\] We divide by 8 : \[n^2+n-2=0\] Then : \[\Delta= 1-4\times(-2)=9\] So : \[n=\frac{1+\sqrt9}{2}=\frac{4}{2}=2\] So : \[a=2n=2\times2=4\]

OpenStudy (anonymous):

but whts the answer

OpenStudy (anonymous):

Ok there is an error

OpenStudy (anonymous):

wht is the answer

OpenStudy (anonymous):

I have to write : \[n=\frac{-1+\sqrt9}{2}=1\] So the smaller number will be : \[a=2\times 1=2\]

OpenStudy (anonymous):

so the answer is 2 right am i correct or no

OpenStudy (anonymous):

Yes, it is 2

OpenStudy (anonymous):

i posted another one